OpenStudy (teller):

Someone mind helping me ? Find the average rate of change of f(x)=-x^2+3x+3 from x=2 to x=6

OpenStudy (inkyvoyd):

(f(6)-f(2))/(6-2)

OpenStudy (teller):

do I divide those?

OpenStudy (inkyvoyd):

f(6)=-15; f(2)=5

OpenStudy (inkyvoyd):

(-15-5)/(6-2)=(-20)/(4)=-5

OpenStudy (teller):

Wait what I' confused .So do I divide what you just posted?

OpenStudy (inkyvoyd):

no. I just worked out the problem. The answer is -5. you want to evaluate the expression \(\Huge \frac{f(b)-f(a)}{(b-a)}\) which in this case is \(\Huge \frac{f(6)-f(2)}{(6-2)}\)

OpenStudy (teller):

Oh okay thank you cause I spent almost ten minutes trying to solve this and the help video was just not working for me .It was confusing me more.

OpenStudy (inkyvoyd):

hahaha no problem :)